OpenStudy (anonymous):

Help I Will Award Metal Write an equation for the line that is parallel to the given line and passes through the given point. y = 5x + 10 (2, 14) A. y = 1^5x + 4 B. y = - 1^5x - 4 C. y = 5x - 68 D. y = 5x + 4

OpenStudy (anonymous):

y-y1=m(x-x1) y-14=5(x-2) y=5x+4 d.

OpenStudy (anonymous):

Thanks!!!!!!